# Asymptotics of the maximum likelihood estimator of the location parameter of Pearson Type VII distribution

### v1

@@ 1. Introduction
$c_{m,7}:=E\left[(\log(1+X_{1}^{2})-F_{m}(0)))^{2}\exp\left(\lambda_{m}\left|\log(1+X_{1}^{2})-F_{m}(0))\right|\right)\right].$
Now let $\lambda:=\delta/c_{m,7}$. ∎
@@
$c_{m,8}:=\dfrac{1}{2}\min\left\{\lambda_{m}(r_{m,2}/3),c_{m,7}\lambda_{m}\right\}.$
Let $\mathcal{B}_{n,1}$ be the event that $\inf_{|t|\geq r_{m,2}/3}L_{n}(t)<F_{m}(0)+c_{m,8}$. Let $\mathcal{B}_{n,2}$ be the event that $L_{n}(0)\geq F_{m}(0)+c_{m,8}$. Then, $\hat{\theta}_{n}\in[-r_{m,2}/2,r_{m,2}/2]$ on the event $\mathcal{B}_{n,1}\cap\mathcal{B}_{n,2}$. Therefore,
@@
$\left|P\left(\hat{\theta}_{n}>\epsilon\right)-P(D_{n}(\epsilon)>0)\right|\leq 2P(\mathcal{A}_{n}^{c})\leq 2c_{m,9}\exp(-c_{m,10}n),\ n\geq 1.$ (5.3)
@@
$H_{m}(\epsilon):=\textup{Var}(D(X_{1},\epsilon))=E\left[D(X_{1},\epsilon)^{2}\right]-G_{m}(\epsilon)^{2}.$
@@
(1) $G_{m}(\epsilon)=-\frac{2m-1}{2(m+1)}\epsilon+O(\epsilon^{2}),\ \epsilon\to+0$.
(2) $H_{m}(\epsilon)=\frac{2m-1}{4m(m+1)}+O(\epsilon),\ \epsilon\to+0$.
@@
$\left|D(X_{1},\epsilon)-D(X_{1},0)-\epsilon\partial_{t}D(X_{1},0)\right|\leq C_{1}\epsilon^{2}.$ (5.4)
@@
$E[D(X_{1},0)]=0,E\left[\partial_{t}D(X_{1},0)\right]=-\frac{2m-1}{2(m+1)}.$
@@
$\left|D(X_{1},\epsilon)^{2}-D(X_{1},0)^{2}-2\epsilon D(X_{1},0)\partial_{t}D(X_{1},0)\right|\leq C_{2}\epsilon^{2}.$ (5.5)
@@
We show (i). We consider the asymptotics of $P(D_{n}(\epsilon)>0)$.
@@
Lemma 5.6 (Petrov [11, Lemma 7.1] [^6] The statement is a little different from [2, Lemma 1]. In [2, Lemma 1], this assertion holds for large $n$, but this is valid for every $n\geq 1$.).
Let $Z_{i},i\geq 1$, be i.i.d. random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}:=\textup{Var}(Z_{1})>0$. Then, for every $n\geq 1$ and every $x\in[0,\sigma^{2}/M]$,
@@
$P(D_{n}(\epsilon)>0)$ $=P\left(\sum_{i=1}^{n}D(X_{i},\epsilon)-G_{m}(\epsilon)>-nG_{m}(\epsilon)\right)$
$\leq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\right).$ (5.6)
@@
$\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\sim\frac{m(2m-1)}{m+1}\epsilon^{2},\ \epsilon\to+0,$ (5.7)
@@
We next give the lower estimate. By Lemma 5.5,
@@
Lemma 5.7 (Petrov [11, Lemma 7.2] [^7] The statement is a little different from [11, Lemma 7.2], however, we can show this assertion in the same manner as in the proof of [11, Lemma 7.2].).
Let $Z_{i},i\geq 1$, be i.i.d. random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}:=\textup{Var}(Z_{1})>0$. Then, for every $\eta>0$, there exists $r>0$ such that for every $x\in[0,r]$, there exists $N$ such that for every $n\geq N$,
@@
$P(D_{n}(\epsilon)>0)\geq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}(1+\eta)\right).$ (5.8)
@@
Now we show (ii), but the proof is almost identical to the proof of (i).
@@
$P(D_{n}(\epsilon/\lambda_{n})>0)\leq\exp\left(-\frac{nG_{m}(\epsilon/\lambda_{n})^{2}}{2H_{m}(\epsilon/\lambda_{n})}\left(1+\frac{G_{m}(\epsilon/\lambda_{n})}{2H_{m}(\epsilon/\lambda_{n})}\right)\right).$
@@
$\lim_{n\to\infty}\lambda_{n}^{2}\frac{G_{m}(\epsilon/\lambda_{n})^{2}}{H_{m}(\epsilon/\lambda_{n})}\left(1+\frac{G_{m}(\epsilon/\lambda_{n})}{2H_{m}(\epsilon/\lambda_{n})}\right)=\frac{m(2m-1)}{m+1}.$
@@
$\limsup_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}\leq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$ (5.9)
@@
$\liminf_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}\geq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$ (5.10)
@@
$\lim_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/\lambda_{n})>0\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
@@
$\lim_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon/\lambda_{n}\right)}{n/\lambda_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
$P\left(\hat{\theta}_{n}<-\epsilon/\lambda_{n}\right)$ can be dealt with in the same manner.
@@
$K\left(\textup{PVII}_{m}(\theta_{1},1)|\textup{PVII}_{m}(\theta_{2},1)\right)=m(F_{m}(\nu_{1}-\nu_{2})-F_{m}(0)).$
@@
$b(\epsilon,\theta):=\inf\left\{K\left(\textup{PVII}_{m}(\theta^{\prime},1)|\textup{PVII}_{m}(\theta,1)\right)\middle||\theta^{\prime}-\theta|>\epsilon\right\}.$
@@
(ii) For the case that $m=1$, the Bahadur efficiency for the joint estimation of the location and the scale is established in [1, Theorem 4] when both the location and the scale are unknown.
@@
We deal with $P\left(\hat{\theta}_{n}>r\right)$. We see that for every $r>0$ and every $\delta>0$,
$P\left(\hat{\theta}_{n}>r\right)\leq P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)+P\left(L_{n}(0)\geq F_{m}(0)+\delta\right).$ (6.1)
We derive upper bounds of these probabilities by modifying the assertions in Section 5. The main difference is $\delta=\delta_{r}$ diverges when $r$ tends to infinity.
First, we give a lemma similar to Lemma 5.2. The proof differs in part. Recall that $\lambda_{m}=\frac{1}{2}(m-\frac{1}{2})$.
@@
$\leq\left(\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]\right)^{n}.$
@@
$E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]$
$E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}},\ X_{1}\geq t/2\right]+E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}},\ X_{1}<t/2\right]$
@@
$\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[(1+(X_{1}-t)^{2}))^{-2\lambda_{m}}\right]=O(t^{-4\lambda_{m}/3}),\ t\to\infty.$
@@
Next, we give a lemma similar to Lemma 5.3. The proof is also similar.
@@
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\frac{F_{m}(r)-F_{m}(0)}{4}\right)\leq c_{m,12}r^{-n\lambda_{m}/2},$
@@
For notational convenience, let $\delta_{r}:=\dfrac{F_{m}(r)-F_{m}(0)}{4}$. Since $|L_{n}^{\prime}(t)|\leq 1$,
$\left\{\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{r}\right\}\subset\bigcup_{k\geq 1}\left\{L_{n}(k\delta_{r}+r)<F_{m}(0)+2\delta_{r}\right\}$
@@
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{r}\right)\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{r}+r)<F_{m}(0)+2\delta_{r}\right)$
$\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{r}+r)<F_{m}(0)+\frac{F_{m}(k\delta_{r}+r)-F_{m}(0)}{2}\right)$
$\leq c_{m,11}\sum_{k=1}^{\infty}(k\delta_{r}+r)^{-n\lambda_{m}}\leq c_{m,11}\delta_{r}\int_{r}^{\infty}x^{-n\lambda_{m}}dx\leq c_{m,11}\delta_{r}r^{1-n\lambda_{m}}.$
By (5.1), $\delta_{r}=O(\log r),r\to\infty$ and we have the assertion. ∎
@@
$P\left(L_{n}(0)\geq F_{m}(0)+\delta_{r}\right)\leq r^{-nc_{m,13}}.$
@@
$P(L_{n}(0)\geq F_{m}(0)+\delta_{r})\leq\left(\exp\left(-\lambda\delta_{r}+\frac{\lambda^{2}}{2}c_{m,7}\right)\right)^{n}.$
By (5.1), there exists a positive constant $r_{m,5}$ such that for every $r>r_{m,5}$, $2c_{m,7}\leq\log r\leq\delta_{r}$. Let $\lambda_{m}^{\prime}:=\min\{1,\lambda_{m}\}$. Thus, for every $r>r_{m,5}$ and every $n\geq 1$,
$P(L_{n}(0)\geq F_{m}(0)+\delta_{r})\leq\left(\exp\left(\left(-\lambda_{m}^{\prime}+\frac{(\lambda_{m}^{\prime})^{2}}{4}\right)\delta_{r}\right)\right)^{n}\leq r^{-nc_{m,13}},$
where we let $c_{m,13}:=\lambda_{m}^{\prime}-\frac{(\lambda_{m}^{\prime})^{2}}{4}>0$. ∎
By applying (6.1) to $\delta=\delta_{r}$, by Lemma 6.2 and Lemma 6.3, there exist positive constants $c_{m},r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
$P\left(\hat{\theta}_{n}>r\right)\leq r^{-c_{m}n}.$
@@
For $M>0$, let $\phi_{M}(x):=x^{2}\wedge M^{2}$. This is bounded and continuous on $\mathbb{R}$. By Theorem 1.2,
$\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}\phi_{M}(x)\varphi_{m}(x)dx,$ (7.1)
where $\varphi_{m}$ is the density function of the distribution $N\left(0,\frac{m+1}{m(2m-1)}\right)$.
@@
$\liminf_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\geq\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx.$ (7.2)
@@
$\limsup_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\leq\frac{m+1}{m(2m-1)}.$ (7.3)
@@
$\lim_{M\to\infty}\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx.$
Hence, it suffices to show that
$\limsup_{M\to\infty}\left(\limsup_{n\to\infty}E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]\right)=0.$ (7.4)
By Fubini’s theorem for non-negative measurable functions and the change of variables $t=\sqrt{s}$,
$E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-M^{2},\ \left|\sqrt{n}\hat{\theta}_{n}\right|\geq M\right]$
$2\int_{M}^{\infty}tP\left(\left|\sqrt{n}\hat{\theta}_{n}\right|>t\right)dt=2n\int_{M/\sqrt{n}}^{\infty}sP\left(\left|\hat{\theta}_{n}\right|>s\right)ds.$
We consider $P\left(\hat{\theta}_{n}>s\right)$.
By (5.7), there exists $\epsilon_{m,3}\in(0,r_{m})$ such that for every $\epsilon\in(0,2\epsilon_{m,3})$,
$\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\geq\frac{m(2m-1)}{4(m+1)}\epsilon^{2}.$ (7.5)
Now we decompose the last integral into three parts:
$\int_{M/\sqrt{n}}^{\infty}=\int_{M/\sqrt{n}}^{\epsilon_{m,3}}+\int_{\epsilon_{m,3}}^{r_{m}+1}+\int_{r_{m}+1}^{\infty},$
@@
$P\left(\hat{\theta}_{n}>s\right)\leq\exp\left(-\frac{m(2m-1)}{4(m+1)}s^{2}n\right)+c_{m,14}\exp(-c_{m,15}n).$
@@
$\limsup_{n\to\infty}2n\int_{M/\sqrt{n}}^{\epsilon_{m,3}}sP\left(\hat{\theta}_{n}>s\right)ds\leq\frac{4(m+1)}{m(2m-1)}\exp\left(-\frac{m(2m-1)}{4(m+1)}M^{2}\right).$ (7.6)
@@
$\limsup_{n\to\infty}2n\int_{\epsilon_{m,3}}^{r_{m}+1}sP\left(\hat{\theta}_{n}>s\right)ds=0.$ (7.7)
@@
$\limsup_{n\to\infty}2n\int_{r_{m}+1}^{\infty}sP\left(\hat{\theta}_{n}>s\right)ds=0.$ (7.8)

### v2

@@ 1. Introduction
$c_{m,7}\coloneqq E\left[(\log(1+X_{1}^{2})-F_{m}(0)))^{2}\exp\left(\lambda_{m}\left|\log(1+X_{1}^{2})-F_{m}(0))\right|\right)\right].$
Now let $\lambda\coloneqq\delta/c_{m,7}$. ∎
@@
$c_{m,8}\coloneqq\dfrac{1}{2}\min\left\{\lambda_{m}(r_{m,2}/3),c_{m,7}\lambda_{m}\right\}.$
Let $\mathcal{B}_{n,1}$ be the event that $\displaystyle\inf_{|t|\geq r_{m,2}/3}L_{n}(t)<F_{m}(0)+c_{m,8}$. Let $\mathcal{B}_{n,2}$ be the event that $L_{n}(0)\geq F_{m}(0)+c_{m,8}$. Then $\hat{\theta}_{n}\in[-r_{m,2}/2,r_{m,2}/2]$ on the event $\mathcal{B}_{n,1}\cap\mathcal{B}_{n,2}$. Therefore,
@@
(5.6) $\left|P\left(\hat{\theta}_{n}>\epsilon\right)-P(D_{n}(\epsilon)>0)\right|\leq 2P(\mathcal{A}_{n}^{c})\leq 2c_{m,9}\exp(-c_{m,10}n),\ n\geq 1.$
@@
$H_{m}(\epsilon)\coloneqq\textup{Var}(D(X_{1},\epsilon))=E\left[D(X_{1},\epsilon)^{2}\right]-G_{m}(\epsilon)^{2}.$
@@
It holds that
(1) $\displaystyle G_{m}(\epsilon)=-\frac{2m-1}{2(m+1)}\epsilon+O(\epsilon^{2}),\ \epsilon\to+0$.
(2) $\displaystyle H_{m}(\epsilon)=\frac{2m-1}{4m(m+1)}+O(\epsilon),\ \epsilon\to+0$.
@@
(5.7) $\left|D(X_{1},\epsilon)-D(X_{1},0)-\epsilon\partial_{t}D(X_{1},0)\right|\leq C_{1}\epsilon^{2}.$
@@
$E\left[D(X_{1},0)\right]=0,E\left[\partial_{t}D(X_{1},0)\right]=-\frac{2m-1}{2(m+1)}.$
@@
(5.8) $\left|D(X_{1},\epsilon)^{2}-D(X_{1},0)^{2}-2\epsilon D(X_{1},0)\partial_{t}D(X_{1},0)\right|\leq C_{2}\epsilon^{2}.$
@@
We show (i) of Theorem 1.4. We consider the asymptotics of $P(D_{n}(\epsilon)>0)$ as $n\to\infty$.
@@
Lemma 5.6 (Petrov [16, Lemma 7.1] [^6] The statement is a little different from [3, Lemma 1]. In [3, Lemma 1], this assertion holds for large $n$, but this is valid for every $n\geq 1$.).
Let $Z_{i},i\geq 1$, be i.i.d. random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}\coloneqq\textup{Var}(Z_{1})>0$. Then, for every $n\geq 1$ and every $x\in[0,\sigma^{2}/M]$,
@@
$\displaystyle P(D_{n}(\epsilon)>0)$ $\displaystyle=P\left(\sum_{i=1}^{n}D(X_{i},\epsilon)-G_{m}(\epsilon)>-nG_{m}(\epsilon)\right)$
(5.9) $\displaystyle\leq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\right).$
@@
(5.10) $\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\sim\frac{m(2m-1)}{m+1}\epsilon^{2},\ \epsilon\to+0,$
@@
We next give the lower estimate (1.2). By Lemma 5.5,
@@
Lemma 5.7 (Petrov [16, Lemma 7.2] [^7] The statement is a little different from [16, Lemma 7.2], however, we can show this assertion in the same manner as in the proof of [16, Lemma 7.2].).
Let $Z_{i},i\geq 1$, be i.i.d. random variables such that $|Z_{1}|\leq M$, $P$-a.s., $E[Z_{1}]=0$, and $\sigma^{2}\coloneqq\textup{Var}(Z_{1})>0$. Then, for every $\eta>0$, there exists $r>0$ such that for every $x\in[0,r]$, there exists $N$ such that for every $n\geq N$,
@@
(5.11) $P(D_{n}(\epsilon)>0)\geq\exp\left(-\frac{nG_{m}(\epsilon)^{2}}{2H_{m}(\epsilon)}(1+\eta)\right).$
@@
The same argument is applicable to $P\left(\hat{\theta}_{n}<-\epsilon\right)$ and we obtain (1.2). Thus the proof of (i) of Theorem 1.4 is completed.
Now we show (ii) of Theorem 1.4, but the proof is almost identical to the proof of (i).
@@
$P(D_{n}(\epsilon/a_{n})>0)\leq\exp\left(-\frac{nG_{m}(\epsilon/a_{n})^{2}}{2H_{m}(\epsilon/a_{n})}\left(1+\frac{G_{m}(\epsilon/a_{n})}{2H_{m}(\epsilon/a_{n})}\right)\right).$
@@
$\lim_{n\to\infty}a_{n}^{2}\frac{G_{m}(\epsilon/a_{n})^{2}}{H_{m}(\epsilon/a_{n})}\left(1+\frac{G_{m}(\epsilon/a_{n})}{2H_{m}(\epsilon/a_{n})}\right)=\frac{m(2m-1)}{m+1}.$
@@
(5.12) $\limsup_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/a_{n})>0\right)}{n/a_{n}^{2}}\leq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
@@
(5.13) $\liminf_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/a_{n})>0\right)}{n/a_{n}^{2}}\geq-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
@@
$\lim_{n\to\infty}\frac{\log P\left(D_{n}(\epsilon/a_{n})>0\right)}{n/a_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
@@
$\lim_{n\to\infty}\frac{\log P\left(\hat{\theta}_{n}>\epsilon/a_{n}\right)}{n/a_{n}^{2}}=-\frac{m(2m-1)}{2(m+1)}\epsilon^{2}.$
$P\left(\hat{\theta}_{n}<-\epsilon/a_{n}\right)$ can be dealt with in the same manner. Thus the proof of (ii) of Theorem 1.4 is completed.
@@
$K\left(\textup{PVII}_{m}(\theta_{1},1)|\textup{PVII}_{m}(\theta_{2},1)\right)=m(F_{m}(\theta_{1}-\theta_{2})-F_{m}(0)).$
@@
$b(\epsilon,\theta)\coloneqq\inf\left\{K\left(\textup{PVII}_{m}(\theta^{\prime},1)|\textup{PVII}_{m}(\theta,1)\right)\middle||\theta^{\prime}-\theta|>\epsilon\right\}.$
@@
(ii) For the case that $m=1$, the Bahadur efficiency for the joint estimation of the location and the scale is established in [2, Theorem 4] when both the location and the scale are unknown. Recently, Akahira [1] showed that the MLE of the location parameter is first order large deviation efficient, which implies the Bahadur efficiency.
@@
We first give an outline of the proof. We estimate the tail probability of $\hat{\theta}_{n}$ by comparing $L_{n}(t)$ at large $|t|$ with its expectation at the true parameter $t=0$, which is $F_{m}(0)$. The probability of the event $\{\hat{\theta}_{n}>r\}$ is controlled by the decomposition (6.1) below. More specifically, if $\hat{\theta}_{n}>r$, then either $L_{n}(t)$ becomes unexpectedly small somewhere on $[r,\infty)$, or $L_{n}(0)$ becomes unexpectedly large.
In order to bound this probability, we control the two terms in (6.1) separately by modifying the assertions in Section 5. Lemma 6.1 below gives a polynomial-type estimate for the lower tail of $L_{n}(t)$ at a fixed large $t$ by the exponential Chebyshev inequality. Then Lemma 6.2 below upgrades this pointwise estimate to the whole tail region $[r,\infty)$ and controls the first term of (6.1), by discretizing the region and using the Lipschitz bound for $L_{n}^{\prime}(t)$. Finally Lemma 6.3 below provides an upper-tail bound for $L_{n}(0)$ and controls the second term of (6.1).
By symmetry, we deal with $P\left(\hat{\theta}_{n}>r\right)$. We see that for every $r>0$ and every $\delta>0$,
(6.1) $P\left(\hat{\theta}_{n}>r\right)\leq P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta\right)+P\left(L_{n}(0)\geq F_{m}(0)+\delta\right).$
First, we give a lemma similar to Lemma 5.2. The proof differs in part. Recall the definition of $\lambda_{m}$ in (5.2).
@@
$\leq\left(\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}}\right]\right)^{n}.$
@@
$E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}}\right]$
$E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}},\ X_{1}\geq t/2\right]+E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}},\ X_{1}<t/2\right]$
@@
$\exp\left(2\lambda_{m}\left(F_{m}(0)+\frac{F_{m}(t)-F_{m}(0)}{2}\right)\right)E\left[\left(1+(X_{1}-t)^{2}\right)^{-2\lambda_{m}}\right]=O(t^{-4\lambda_{m}/3}),\ t\to\infty.$
@@
Next, we give a lemma similar to Lemma 5.3. The proof is also similar. Recall the definition of $\delta_{m}(r)$ in (5.3).
@@
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{m}(r)\right)\leq c_{m,12}r^{-n\lambda_{m}/2},$
@@
Since $|L_{n}^{\prime}(t)|\leq 1$,
$\left\{\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{m}(r)\right\}\subset\bigcup_{k\geq 1}\left\{L_{n}(k\delta_{m}(r)+r)<F_{m}(0)+2\delta_{m}(r)\right\}$
@@
$P\left(\inf_{t\geq r}L_{n}(t)<F_{m}(0)+\delta_{m}(r)\right)\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{m}(r)+r)<F_{m}(0)+2\delta_{m}(r)\right)$
$\leq\sum_{k=1}^{\infty}P\left(L_{n}(k\delta_{m}(r)+r)<F_{m}(0)+\frac{F_{m}(k\delta_{m}(r)+r)-F_{m}(0)}{2}\right)$
$\leq c_{m,11}\sum_{k=1}^{\infty}(k\delta_{m}(r)+r)^{-n\lambda_{m}}\leq c_{m,11}\delta_{m}(r)\int_{r}^{\infty}x^{-n\lambda_{m}}dx\leq c_{m,11}\delta_{m}(r)r^{1-n\lambda_{m}}.$
By (5.1), $\delta_{m}(r)=O(\log r),r\to\infty$ and we have the assertion. ∎
@@
$P\left(L_{n}(0)\geq F_{m}(0)+\delta_{m}(r)\right)\leq r^{-nc_{m,13}}.$
@@
$P(L_{n}(0)\geq F_{m}(0)+\delta_{m}(r))\leq\left(\exp\left(-\lambda\delta_{m}(r)+\frac{\lambda^{2}}{2}c_{m,7}\right)\right)^{n}.$
By (5.1), there exists a positive constant $r_{m,5}$ such that for every $r>r_{m,5}$, $2c_{m,7}\leq\log r\leq\delta_{m}(r)$. Let $\lambda_{m}^{\prime}\coloneqq\min\{1,\lambda_{m}\}$. Thus, for every $r>r_{m,5}$ and every $n\geq 1$,
$P(L_{n}(0)\geq F_{m}(0)+\delta_{m}(r))\leq\left(\exp\left(\left(-\lambda_{m}^{\prime}+\frac{(\lambda_{m}^{\prime})^{2}}{4}\right)\delta_{m}(r)\right)\right)^{n}\leq r^{-nc_{m,13}},$
where we let $c_{m,13}\coloneqq\lambda_{m}^{\prime}-\dfrac{(\lambda_{m}^{\prime})^{2}}{4}>0$. ∎
By applying (6.1) to $\delta=\delta_{m}(r)$, by Lemma 6.2 and Lemma 6.3, there exist positive constants $r_{m}$ and $N_{m}\in\mathbb{N}$ depending only on $m$ such that for every $r\geq r_{m}$ and every $n\geq N_{m}$,
$P\left(\hat{\theta}_{n}>r\right)\leq r^{-c_{m,13}n}.$
@@
We first give an outline of the proof. The proof is a uniform-integrability argument based on the asymptotic normality (Theorem 1.2) and the Bahadur efficiency (two estimates used in the proof of Theorem 1.4) and the tail bound (Theorem 1.5). We show the lower and upper bounds separately. The lower bound is an easy consequence of Theorem 1.2.
In order to obtain the upper bound, it suffices to show that the family $\left\{\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}\right\}_{n}$ is uniformly integrable, which is reduced to (7.4) below. The key step is to rewrite the tail contribution as an integral of tail probabilities as in (7) and then split the integral into three regimes as in (7.7). For the small-to-moderate region and the intermediate region, we use (5.6) and (5) used in the proof of Theorem 1.4 respectively. For the far tail region, we apply Theorem 1.5.
For $M>0$, let $\phi_{M}(x)\coloneqq\min\{x^{2},M^{2}\}$. This is bounded and continuous on $\mathbb{R}$. Recall that $\varphi_{m}$ is the density function of the distribution $\displaystyle N\left(0,\dfrac{m+1}{m(2m-1)}\right)$. By Theorem 1.2,
(7.1) $\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}\phi_{M}(x)\varphi_{m}(x)dx.$
@@
(7.2) $\liminf_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\geq\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx=\frac{m+1}{m(2m-1)}.$
@@
(7.3) $\limsup_{n\to\infty}nE\left[\left(\hat{\theta}_{n}\right)^{2}\right]\leq\frac{m+1}{m(2m-1)}.$
@@
$\lim_{M\to\infty}\lim_{n\to\infty}E\left[\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]=\int_{\mathbb{R}}x^{2}\varphi_{m}(x)dx=\frac{m+1}{m(2m-1)}.$
Hence it suffices to show that
(7.4) $\limsup_{M\to\infty}\left(\limsup_{n\to\infty}E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]\right)=0.$
By Fubini’s theorem for non-negative measurable functions and the change of variables $t=\sqrt{s}$, we obtain that
$\displaystyle E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-\phi_{M}\left(\sqrt{n}\hat{\theta}_{n}\right)\right]$ $\displaystyle=E\left[\left(\sqrt{n}\hat{\theta}_{n}\right)^{2}-M^{2},\ \left|\sqrt{n}\hat{\theta}_{n}\right|\geq M\right]$
$\displaystyle=2\int_{M}^{\infty}tP\left(\left|\sqrt{n}\hat{\theta}_{n}\right|>t\right)dt$
(7.5) $\displaystyle=2n\int_{M/\sqrt{n}}^{\infty}sP\left(\left|\hat{\theta}_{n}\right|>s\right)ds.$
By symmetry, we consider $P\left(\hat{\theta}_{n}>s\right)$. By (5.10), there exists $\epsilon_{m,3}\in(0,r_{m})$ such that for every $\epsilon\in(0,2\epsilon_{m,3})$,
(7.6) $\frac{G_{m}(\epsilon)^{2}}{H_{m}(\epsilon)}\left(1+\frac{G_{m}(\epsilon)}{2H_{m}(\epsilon)}\right)\geq\frac{m(2m-1)}{4(m+1)}\epsilon^{2}.$
Now we decompose the last integral in (7) into three parts:
(7.7) $\int_{M/\sqrt{n}}^{\infty}=\int_{M/\sqrt{n}}^{\epsilon_{m,3}}+\int_{\epsilon_{m,3}}^{r_{m}+1}+\int_{r_{m}+1}^{\infty},$
@@
(7.8) $P\left(\hat{\theta}_{n}>s\right)\leq\exp\left(-\frac{m(2m-1)}{4(m+1)}s^{2}n\right)+c_{m,14}\exp(-c_{m,15}n).$
@@
(7.9) $\limsup_{n\to\infty}2n\int_{M/\sqrt{n}}^{\epsilon_{m,3}}sP\left(\hat{\theta}_{n}>s\right)ds\leq\frac{4(m+1)}{m(2m-1)}\exp\left(-\frac{m(2m-1)}{4(m+1)}M^{2}\right).$
@@
by applying (7.8) to $s=\epsilon_{m,3}$,
(7.10) $\limsup_{n\to\infty}2n\int_{\epsilon_{m,3}}^{r_{m}+1}sP\left(\hat{\theta}_{n}>s\right)ds=0.$
@@
(7.11) $\limsup_{n\to\infty}2n\int_{r_{m}+1}^{\infty}sP\left(\hat{\theta}_{n}>s\right)ds=0.$